Write the equation of a line that is parallel to 3x+y=10 and runs through (9,10)
Answer by stanbon(75887)

Write the equation of a line that is
parallel to 3x+y=10 and runs through (9,10)
----
The given line has slope = -3
----
Any line parallel to it must have a slope of -3.
---
Equation Form:
y = mx + b
10 = -3*9 + b
b = 37
----
Equation:
y = -3x+37
===========
Cheers,
Stan H.
===========